OpenStudy (anonymous):

please help!!!!!!!! f(x)=3x-6 g(x)=x-2 find f(g) and domain

OpenStudy (anonymous):

Basically you need to put in G(x) into all the x values of the F(x) function. f(g) = 3(x-2) - 6 f(g) = 3x-6-6 f(g) = 3x-12 --> final answer.

OpenStudy (marigirl):

You can think of it like this: f(x)=3x-6 and rewrite it to f(x)= 3*unit -6 g(x)=x-2 and rewrite it to g(x)= unit-2 find f(g) and domain f(g)= f(x-2) now we know f(x) is 3*unit -6 and our unit now is (x-2) f(g)= 3(x-2) - 6 f(g)= 3x-6-6 Expand brackets f(g)= 3x-12 Domain can be real numbers since both f(g) and g(x) are functions with real values

OpenStudy (anonymous):

@marigirl @kieran01pd2016 i know it equals 3x-12 but the answers only supplies me with a)3 b)-3 c)1

OpenStudy (anonymous):

fg has to be a single number for the answer i just dont know what

jimthompson5910 (jim_thompson5910):

`f(g)` makes no sense to be honest it should be `f(g(x))` or something like `f(g(23))`

OpenStudy (dinamix):

its -3

jimthompson5910 (jim_thompson5910):

I guess if you did say `f(g)`, it would mean you replace every x in f(x) with g f(x)=3x-6 f(g)=3g-6 but that's not what they want

OpenStudy (irishboy123):

scan or link the question, pls :p

OpenStudy (anonymous):

@irishboy123 sadly my class supplied this exact question only with nothing else

OpenStudy (anonymous):

It's 3 guys thanks .

OpenStudy (anonymous):

I think idk just gonna guess thanks

OpenStudy (anonymous):

Well, good luck yo!

OpenStudy (anonymous):

Let f(x) = 3x - 6 and g(x) = x - 2. Find (f*g) and states its domain? A. 3; all real numbers except x = 2 B. –3; all real numbers except x = 3 C. 1; all real numbers D. 3; all real numbers

jimthompson5910 (jim_thompson5910):

did you mean (f/g)(x) ?

jimthompson5910 (jim_thompson5910):

or f(x)/g(x) ?

OpenStudy (anonymous):

It's f times g , it makes no sense to me but it's the question

jimthompson5910 (jim_thompson5910):

This is a very awful question. I think they meant to ask f/g because none of the answer choices make sense if it was f*g

jimthompson5910 (jim_thompson5910):

\[\Large (f/g)(x) = \frac{f(x)}{g(x)}\] \[\Large (f/g)(x) = \frac{3x-6}{x-2}\] \[\Large (f/g)(x) = \frac{3(x-2)}{x-2}\] \[\Large (f/g)(x) = \frac{3\cancel{(x-2)}}{\cancel{x-2}}\] \[\Large (f/g)(x) = 3\] where x cannot equal 2 (to avoid division by zero errors)

OpenStudy (anonymous):

yeah ill ask about if it was f/g would the answer be a

jimthompson5910 (jim_thompson5910):

IF it was f*g, then you'd have \[\Large (f*g)(x) = f(x)*g(x)\] \[\Large (f*g)(x) = (3x-6)*(x-2)\] \[\Large (f*g)(x) = 3x^2-6x-6x+12\] \[\Large (f*g)(x) = 3x^2-12x+12\]